Cost Flow Methods

The following three identical units of Item LO3V are purchased during April:

Item Beta Units Cost

April 2Purchase 1 $144

April 15Purchase 1 145

April 20Purchase 1 146

Total  3 $435

Average cost per unit    $145($435 ÷ 3 units)

Assume that one unit is sold on April 27 for $190. Determine the gross profit for April and ending inventory on April 30 using the (a) first-in, first-out (FIFO); (b) last-in, first-out (LIFO); and (c) weighted average cost method.

Gross ProfitEnding Inventory

a. First-in, first-out (FIFO)$$

b. Last-in, first-out (LIFO)$$

c. Weighted average cost$$
Answer #1

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet. For detailed answer refer to the supporting sheet.

Answer Gross Profit as per FIFO = sales - Cost of good sold as per FIFO = 190-144 = $ 46 Cost of ending inventory FIFO = No.

Gross Profit as Per Average cost = sales - Cost of good sold as per Average cost + = 190 -(1*145) = $ 45 Cost of ending inven